WATER CLUSTERS OBSERVED BY CHIRPED-PULSE ROTATIONAL SPECTROSCOPY: STRUCTURES AND HYDROGEN BONDING

\underline{ZBIGNIEW KISIEL}, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warszawa, Poland; CRISTOBAL PEREZ, MATT T. MUCKLE, DANIEL P. ZALESKI, NATHAN SEIFERT, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904-4319; BERHANE TEMELSO, GEORGE C. SHIELDS, Dean’s Office, College of Arts and Sciences, and Department of Chemistry, Bucknell University, Lewisburg, PA 17837.

The chirped-pulse rotational spectra of the hexamer, heptamer, and nonamer water clusters have been assigned and analysed as described in the preceding talk. The observation of all singly substituted ^{18}O isotopologues for several clusters allowed direct determination of the r_{s} geometries of their oxygen atom frameworks. The r_{s} analysis was then complemented by least-squares determination of the r_{0} geometries, which circumvented the problems with some imaginary substitution coordinates and allowed direct comparisons with computations. For the cage, prism and the book water hexamer clusters the agreement in O⋯O distances between experiment and vibrationally averaged computed geometries is typically at the 0.01Å level. This is a previously unanticipated level of agreement between experiment and theory, which facilitates more confident discussion of the internal properties of these clusters. The hexamer and larger water clusters begin to display the diversity of hydrogen bonding that is characteristic of condensed water and various aspects of this behaviour are discussed.